OpenStudy (anonymous):

2013 + k is a perfect square . Find all the possible values of k.

OpenStudy (mr.math):

You want to find all values of k such that \(2013+k=n^2 \implies k=n^2-2013\), which has infinitely many solutions over the integers. Thus there are infinitely many k such that \(2013+k\) is a perfect square.

OpenStudy (dumbcow):

x^2 = 2013 + k -> k = x^2 -2013, for all integers x>0

OpenStudy (anonymous):

k>=-2013

OpenStudy (anonymous):

and the smallest k is =-2012

OpenStudy (anonymous):

k=-2013 is that possible?

OpenStudy (anonymous):

(a+b+c)^2 = (2013 + k) is the original equation.

OpenStudy (anonymous):

are we looking for solution in R or C?

OpenStudy (anonymous):

In integers.

OpenStudy (anonymous):

Mr. Math and dumbcow seem right..

OpenStudy (anonymous):

a,b,c are also integers?

OpenStudy (anonymous):

yes

OpenStudy (earthcitizen):

there should be only two values of k ?

OpenStudy (anonymous):

no,you are supposed to find all the possible values .

OpenStudy (anonymous):

all possible values seem infinite..

OpenStudy (anonymous):

1 more condition : a^ + 2bc = 2012, b^2 + 2ca = 1,c^2 + 2ab = k.

OpenStudy (anonymous):

(:

OpenStudy (mr.math):

Lol, Could your write the whole problem so we can help?

OpenStudy (anonymous):

If a, b, c are integers such that a^ + 2bc = 2012, b^2 + 2ca = 1,c^2 + 2ab = k.Then,find all the possible values of k.

OpenStudy (anonymous):

\[ac \le0 \qquad a \le0 \quad or \quad c \le0\]

OpenStudy (anonymous):

is there any options?